can someone help me plss before monday?
solve the equation.
20. a+3/a = 5

Answer by stanbon(75887)

a+3/a = 5
---
There are two ways to interpret your posting:
---
1st:
a + (3/a) = 5
a^2 + 3 = 5a
a^2-5a+3 = 0
Then use the quadratic formula to solve for "a":
--------------------------
2nd:
(a+3)/a = 5
Cross-multiply to get:
a+3 = 5a
4a = 3
a = 3/4
=================
Cheers,
Stan H.
=====================